A load W = 6000N is supported by two ropes one of which makes an angle of

     50o with the horizontal while the other rope makes an angle of 60o with the

     Horizontal. Calculate the force exerted by each rope so as to provide

      equilibrium. CALCULATE the Resultant, and Resultant Angle.



1
Expert's answer
2021-09-04T07:35:07-0400
"\sum F_x=0:\quad T_1\cos50^{\circ}=T_2\cos60^{\circ}"

"T_1=0.778T_2"

"\sum F_y=0:\quad T_1\sin50^{\circ}+T_2\sin60^{\circ}=W"

"0.778T_2+0.866T_2=6000"

"T_2=3628\:{\rm N}, \quad T_1=2822\:\rm N"